# Gradient-Based Multi-Objective Deep Learning: Algorithms, Theories, Applications, and Beyond

### v1

@@ 5.1 Network Structure
(a) Hypernetwork-based methods.
(b) Preference-conditioned network-based methods.
(c) Model combination-based methods.
@@ 5.1.1 Hypernetwork
Hypernetwork is a neural network that generates the parameters for another target network [85]. This idea has been leveraged by the Pareto Hypernetwork (PHN) [43] and Controllable Pareto Multi-Task Learning (CPMTL) [44] to learn the entire Pareto set, where the input is the user preference ${\bm{\alpha}}$ and the output is the target network’s parameters. The hypernetwork usually consists of several MLP layers, and this framework has been adopted in many subsequent approaches such as PHN-HVI [45] and SUHNPF [46]. Recently, Tuan et al. [47] propose using a transformer architecture as the hypernetwork, which demonstrates superior performance compared to MLP-based hypernetwork models.
A primary limitation of hypernetwork-based algorithms is their size. Since the output dimension matches the number of parameters in the target networks, hypernetworks are often much larger than the base networks, limiting their applicability to large models. Some methods address this limitation by employing hypernetworks with chunking [43, 44]. Chunking involves partitioning the parameter space into smaller, more manageable segments, enabling the hypernetwork to generate parameters more efficiently and scalability. This approach reduces the overall size of the hypernetwork while preserving its capability to produce accurate parameters for the target network.
@@ 5.1.2 Preference-Conditioned Network
In preference-conditioned network-based methods, instead of using a hypernetwork to generate weights, the original model is directly modified to incorporate preferences. COSMOS [48] suggests adding preferences as an additional input to the model by combining user preferences ${\bm{\alpha}}$ with input data ${\bm{x}}$, thereby increasing the input dimension of the original neural network.
However, this input-based conditioning has limited ability to produce diverse solutions. Some studies [49, 34, 50] employ the Feature-wise Linear Modulation (FiLM) layer [86] to condition the network. The FiLM layer works by applying an affine transformation to the feature maps. Specifically, given a feature map ${\bm{u}}$ with $C$ channels, we use an MLP to generate conditioning parameters ${\bm{\gamma}}\in\mathbb{R}^{C}$ and ${\bm{\beta}}\in\mathbb{R}^{C}$ based on the given preference ${\bm{\alpha}}$. Then, the FiLM layer modifies the features as follows:
${\bm{u}}^{\prime}_{c}=\gamma_{c}\cdot{\bm{u}}_{c}+\beta_{c},$ (39)
where ${\bm{u}}_{c}$ represents the $c$-th channel of ${\bm{u}}$. This method achieves better conditioning ability compared to merely adding the preference to the input layer. Raychaudhuri et al. [50] also propose leveraging another network to predict the network architecture. This approach dynamically adapts the model’s structure according to user preferences, allowing for more flexible and efficient learning.
@@ 5.1.3 Model Combination
Model combination-based methods construct a composite model by integrating multiple individual models, thereby offering an effective way to introduce diversity. PaMaL [51] achieves this by learning several base models and combining them through a weighted sum of their parameters based on user preferences:
@@
Although this method incorporates more parameters than preference-conditioned network-based methods, PaMaL enhances diversity, providing users with a broader range of choices.
Despite its benefits, learning multiple models simultaneously may become inefficient when handling a large number of objectives. To mitigate this issue, LORPMAN [52] proposes learning a full-rank based model, ${\bm{\theta}}_{0}$, and several low-rank models, ${\bm{B}}_{i}{\bm{A}}_{i}$, where ${\bm{B}}_{i}\in\mathbb{R}^{d\times r}$ and ${\bm{A}}_{i}\in\mathbb{R}^{r\times k}$. Given a user preference ${\bm{\alpha}}$, the composite model is given by:
@@
where ${\bm{C}}\in\mathbb{R}^{r\times r\times m}$ represents a learnable core tensor, ${\bm{A}}\in\mathbb{R}^{r\times d}$, ${\bm{B}}\in\mathbb{R}^{r\times k}$ are learnable matrices, and ×$\times_{u}$ represents the mode-$u$ tensor multiplication [87].
Furthermore, Zhong et al.[54] address scalability challenges for large-scale models, such as language models, using SVD decomposition:
@@
where ${\bm{U}}\in\mathbb{R}^{d\times(r+m)}$, ${\bm{V}}\in\mathbb{R}^{(r+m)\times k}$, and ${\bm{\Sigma}}$ is a diagonal matrix defined as $\text{diag}(\sigma_{1},\cdots,\sigma_{r},s\alpha_{1},\cdots,s\alpha_{m})$.
Instead of training new models, Tang et al. [55] propose to utilize a mixture of experts [88] method to combine pre-existing models in order to achieve the Pareto set. Given $n$ models, ${\bm{\theta}}_{1},\cdots,{\bm{\theta}}_{n}$, each trained on different datasets, an initial step involves integrating these models into a unified base model, ${\bm{\theta}}_{0}$. For some modules in the network, differences between the individual models and the base model, denoted as ${\bm{\theta}}_{i}-{\bm{\theta}}_{0}$, are maintained. These differences are treated as experts within the framework. Then, the mixture of experts method is applied, where a gating network, conditioned on user preferences, determines the weighting of these experts. It enables the generation of diverse models by applying varying weighted combinations of experts to align with specific user preferences.
@@ 5.2 Training Strategy
This section outlines the training approach for the parameters of the network structures introduced in Section 5.1, referred to as ${\bm{\phi}}$. Specifically, ${\bm{\phi}}$ refer to those of the Hypernetwork discussed in Section 5.1.1, the parameters of the preference-conditioned network described in Section 5.1.2, or the parameters of the base models or low-rank models presented in Section 5.1.3. Current algorithms follow similar approaches, with the core idea being to sample user preferences and optimize the structure parameters to generate networks that align with those preferences. The main idea is to sample user preferences and optimize the structure parameters to produce networks that align with these preferences. The general training objective can be written as:
@@
where $({\bm{x}},{\bm{y}})$ are training data sampled from dataset $\mathcal{D}$, ${\bm{\ell}}(\cdot)$ is the multi-objective loss function and $\tilde{g}_{\bm{\alpha}}(\cdot)$ is an MOO algorithm that produces a single solution given preference vector ${\bm{\alpha}}$. In general, most optimization algorithms discussed in Sections 2.2 and 4 can be adopted. Below, we summarize the algorithms adopted by existing methods.
@@
  • Preference-Aware Weighting Methods: As discussed in Section 4.1, these methods incorporate user preferences into optimization. PHN [43] employs the Exact Pareto Optimal (EPO) solver [29] to find Pareto optimal solutions aligned with preferences. CPMTL [44] uses a constrained approach inspired by PMTL [28].
  • Hypervolume Maximization: As discussed in Section 4.2, hypervolume maximization can simultaneously optimize both the diversity and convergence of a solution set. PHN-HVI [45] leverages hypervolume maximization to optimize the structural parameters ${\bm{\phi}}$. In each iteration, it samples a set of preference vectors and generates solutions based on these vectors. Subsequently, hypervolume maximization is applied to promote both the diversity and convergence of the solutions.
@@ 5.3 Discussions
Selecting the appropriate structural framework and optimization algorithms is important for effectively learning an infinite Pareto set.
Hypernetwork-based approaches offer significant flexibility by generating a diverse set of models tailored to varying user preferences. However, a primary limitation of hypernetworks is their size. Since the output dimension directly corresponds to the number of parameters in the target network, hypernetworks can become substantially larger than the base models. This increase in size can limit their applicability, particularly for large-scale networks where computational and memory resources are constrained.
In contrast, preference-conditioned network-based methods are highly parameter-efficient. They introduce only a minimal number of additional parameters relative to the base model, making them well-suited for large-scale applications where adding extra parameters is costly. Despite their efficiency, these methods may struggle to provide sufficient diversity within the parameter space. This limitation can hinder their ability to approximate complex PS accurately, as the restricted capacity may fail to capture all the nuances of Pareto optimal solutions.
Model combination-based approaches serve as a middle ground between hypernetwork and preference-conditioned network methods. By integrating multiple models, they can offer a robust approximation of the PS while maintaining scalability for large models and handling multiple objectives effectively. These methods allow for parameter adjustments through the selection of appropriate ranks in low-rank approximations.
Regarding training strategies, most algorithms use linear scalarization for its simplicity and strong performance. However, when aligning solutions with user preferences is important, preference-dependent weighting schemes discussed in Section 4.1 are more suitable. Additionally, to achieve HV optimal solutions, HV-based optimization criteria discussed in Section 4.2 can also be employed, as they better capture the PF by focusing on the dominated objective space volume.
@@ 6 Theories
This section reviews the paper that provides a theoretical analysis of the convergence of gradient-balancing methods in Section 3.2. The convergence property is first examined under the deterministic gradient setting (i.e., full-batch gradient descent). Subsequently, the convergence property is analyzed under the stochastic gradient setting (i.e., stochastic gradient descent). Note that in non-convex MOO, convergence refers to reaching Pareto stationary, as defined in Definition 6.
@@ 6.1.1 Deterministic Gradient
When dealing with deterministic gradients, it has shown that under mild conditions, MGDA [6] can converge to a Pareto stationary point with a convergence rate of the order -1/2 [89], where $K$ is the number of iterations. This rate is comparable to that of single-objective optimization [90]. The fundamental idea is to assess the reduction in each individual objective function. When the current solution is not Pareto stationary, it can be shown that the common descent direction ${\bm{d}}$ identified by MGDA ensures an improvement of at least $\mathcal{O}(\left\|{\bm{d}}\right\|^{2})$ in each objective, which mirrors the situation in single-objective optimization. Therefore, the method used in single-objective optimization can be applied to prove the convergence. Furthermore, other algorithms such as CAGrad [19], Nash-MTL [21], and PCGrad [25] also offer convergence analyses in similar approaches.
TABLE II: Convergence of existing stochastic gradient MOO algorithms. LS: $L$-smooth, BG: bounded gradient, BF: bounded function value, GS: generalized $L$-smooth, CA: conflict-avoidant.
Method Batch Size Assumption Sample Complexity Bounded CA Distance
@@ 6.1.2 Stochastic Gradient
Liu and Vicente [91] are the first to analyze the convergence of gradient-based MOO algorithms under stochastic gradients. They propose a stochastic version of MGDA [6] that computes a common descent direction using stochastic gradients for all objectives and prove convergence to Pareto optimal solutions under the assumption of convex objective functions. However, due to the inherent bias of common descent direction ${\bm{d}}$ introduced by stochastic gradient estimations, their analysis requires the use of an increasing batch size that grows linearly with the number of iterations to ensure convergence. This requirement can be impractical in real-world applications, as large batch sizes lead to increased computational costs and memory usage.
To overcome the limitations associated with increasing batch sizes, Zhou et al. [92] propose the Correlation-Reduced Stochastic Multi-Objective Gradient Manipulation (CR-MOGM) method. It addresses the bias in the common descent direction by introducing a smoothing technique on weight coefficient ${\bm{\lambda}}$. In CR-MOGM, the weight coefficient ${\bm{\lambda}}^{(k)}$ at iteration $k$ is updated using a moving average:
@@
where $\gamma\in[0,1]$ is a smoothing factor, $\hat{{\bm{\lambda}}}^{(k)}$ is the weight vector obtained by the MOO solver (such as MGDA) at iteration $k$ using the current stochastic gradients, and ${\bm{\lambda}}^{(k-1)}$ is the smoothed weight from the previous iteration. This smoothing reduces the variance in the weight, leading to a more stable and reliable common descent direction. Xu et al. [94] further analyze the convergence by updating the objective weights every $\tau$ iterations instead of at each iteration and achieve the same convergence rate ${\mathcal{O}}(\epsilon^{-2})$ as CR-MOGM. They also match the convergence rate in stochastic single-objective optimization [97].
@@
$\hat{{\bm{g}}}^{(k+1)}_{i}=\prod_{L_{i}}\left(\hat{{\bm{g}}}^{(k)}_{i}-\gamma\hat{{\bm{g}}}^{(k)}_{i}-{\bm{g}}^{(k)}_{i}\right),$ (46)
where $\prod_{L_{i}}$ denotes the projection to set $\{{\bm{g}}\in\mathbb{R}^{d}~{}~{}|~{}~{}\left\|{\bm{g}}\right\|\leq L_{i}\}$, and $L_{i}$ is the Lipschitz constant of $f_{i}({\bm{\theta}})$. However, the analysis requires the number of iterations $K$ to be sufficiently large (on the order of ${\mathcal{O}}(m^{10})$). This high dependency on $m$ makes MoCo less practical for problems involving many objectives. Additionally, the convergence analysis of CR-MOGM and MoCo relies on the assumption that the objective functions have bounded values.
To address these limitations, Chen et al. [95] propose the Multi-objective gradient with double sampling algorithm (MoDo). MoDo mitigates bias in stochastic gradient-based MOO methods without requiring increasing batch sizes or assuming bounded function values. It introduces a double sampling technique to obtain unbiased estimates of the gradient products needed for updating the weight coefficients. Specifically, at each iteration, MoDo updates the weight vector ${\bm{\lambda}}^{(k)}$ using gradients computed on two independent mini-batches:
@@
where $\eta$ is the step size, $\prod_{\Delta_{m-1}}$ denotes projection onto the simplex $\Delta_{m-1}$, and ${\bm{G}}^{(k)}({{\bm{z}}^{(k)}_{1}})$, ${\bm{G}}^{(k)}({{\bm{z}}^{(k)}_{2}})$ are the stochastic gradients evaluated on two independent mini-batches ${\bm{z}}_{1}^{(k)}$ and ${\bm{z}}_{2}^{(k)}$. By reducing the bias in estimating the common descent direction, MoDo achieves convergence guarantees without bounded function value assumption. Additionally, it guarantees a bounded conflict-avoidant (CA) distance, which is the distance between the estimated update direction ${\bm{d}}$ and the conflict-avoidant direction defined in problem (12).
Similarly, Xiao et al. [20] propose the Stochastic Direction-oriented Multi-objective Gradient descent (SDMGrad) algorithm, which introduces a new direction-oriented multi-objective formulation by regularizing the common descent direction within a neighborhood of a target direction, such as the average gradient of all objectives. In SDMGrad, the weight vector ${\bm{\lambda}}^{(k)}$ is updated by:
@@
where $\eta$ is the step size, $\gamma$ is a regularization factor controlling the proximity to the target direction ${\bm{g}}_{0}({\bm{z}}_{2}^{(k)})$ (e.g., the weighted sum of gradients). By incorporating this regularization, SDMGrad effectively balances optimization across objectives while guiding the overall descent direction. Zhang et al. [96] further consider the convergence under generalized $L$-smoothness [98] and without bounded gradient assumption. A summary of convergence of existing stochastic gradient MOO algorithms is provided in Table II.

### v2

@@ 5.1 Network Structure
(a) Methods based on the hypernetwork.
(b) Methods based on preference-conditioned network.
(c) Methods based on model combination.
@@ 5.1.1 Hypernetwork
Hypernetwork is a neural network that generates the parameters for another target network [93]. This idea has been leveraged by the Pareto Hypernetwork (PHN) [52] and Controllable Pareto Multi-Task Learning (CPMTL) [53] to learn the entire Pareto set, where the input is the user preference ${\bm{\alpha}}$ and the output is the target network’s parameters. The hypernetwork usually consists of several MLP layers, and has been adopted in many subsequent approaches such as PHN-HVI [54] and SUHNPF [55]. Recently, Tuan et al. [56] propose using a transformer architecture as the hypernetwork, which demonstrates superior performance compared to MLP-based hypernetwork models.
A primary limitation of hypernetwork-based algorithms is their sizes. Since the output dimension matches the number of parameters in the target network, hypernetworks are often much larger than the base networks, limiting their applicability to large models. Some methods address this limitation by employing hypernetworks with chunking [52, 53]. Chunking involves partitioning the parameter space into smaller, more manageable segments, enabling the hypernetwork to generate parameters more efficiently and scalability. This approach reduces the overall size of the hypernetwork while preserving its capability to produce accurate parameters for the target network.
@@ 5.1.2 Preference-Conditioned Network
Instead of using a hypernetwork to generate weights, the original model can be directly modified to incorporate preferences. COSMOS [57] suggests adding preferences as an additional input to the model by combining user preference ${\bm{\alpha}}$ with input data ${\bm{x}}$, thereby increasing the input dimension of the original neural network.
However, such input-based conditioning has limited ability to produce diverse solutions. Some studies [58, 42, 59] employ the Feature-wise Linear Modulation (FiLM) layer [94] to condition the network. The FiLM layer works by applying an affine transformation to the feature maps. Specifically, given a feature map ${\bm{u}}$ with $C$ channels, we use an MLP to generate conditioning parameters ${\bm{\gamma}}\in\mathbb{R}^{C}$ and ${\bm{\beta}}\in\mathbb{R}^{C}$ based on the given preference ${\bm{\alpha}}$. Then, the FiLM layer modifies the features as ${\bm{u}}^{\prime}_{c}=\gamma_{c}\cdot{\bm{u}}_{c}+\beta_{c},$ where ${\bm{u}}_{c}$ is the $c$-th channel of ${\bm{u}}$. This method achieves better conditioning ability compared to merely adding the preference to the input layer. Raychaudhuri et al. [59] also propose leveraging another network to predict the network architecture. It dynamically adapts the model’s structure according to user preferences, allowing for more flexible and efficient learning.
@@ 5.1.3 Model Combination
Methods based on model combination construct a composite model by integrating multiple individual models, thereby offering an effective way to introduce diversity. PaMaL [60] achieves this by learning several base models and combining them through a weighted sum of their parameters based on user preferences:
@@
Although this incorporates more parameters than methods based on preference-conditioned networks, PaMaL enhances diversity and provides users with a broader range of choices.
Despite its benefits, learning multiple models simultaneously may become inefficient when handling a large number of objectives. To mitigate this issue, LORPMAN [61] proposes learning a full-rank model ${\bm{\theta}}_{0}\in\mathbb{R}^{p\times q}$ and $m$ low-rank models ${\bm{B}}_{i}{\bm{A}}_{i}$’s, where ${\bm{B}}_{i}\in\mathbb{R}^{p\times r}$ and ${\bm{A}}_{i}\in\mathbb{R}^{r\times q}$. Given user preference ${\bm{\alpha}}$, the composite model is given by:
@@
where ${\bm{C}}\in\mathbb{R}^{r\times r\times m}$ is a learnable core tensor, ${\bm{A}}\in\mathbb{R}^{r\times p}$, ${\bm{B}}\in\mathbb{R}^{r\times q}$ are learnable matrices, and ×$\times_{u}$ denotes mode-$u$ tensor multiplication [95]. Zhong et al.[63] address scalability challenges for large-scale models, such as language models, using SVD:
@@
where ${\bm{U}}\in\mathbb{R}^{p\times(r+m)}$ and ${\bm{V}}\in\mathbb{R}^{(r+m)\times q}$ are learnable matrices. ${\bm{\Sigma}}$ is a diagonal matrix defined as $\text{diag}(\sigma_{1},\dots,\sigma_{r},s\alpha_{1},\dots,s\alpha_{m})$, where $\{\sigma_{i}\}_{i=1}^{r}$ and $s$ are learnable scalars.
Instead of training new models, Tang et al. [64] propose to utilize a mixture of experts [96] to combine pre-existing models in order to achieve the Pareto set. Given $n$ models, ${\bm{\theta}}_{1},\dots,{\bm{\theta}}_{n}$, each trained on different datasets, they are first integrated to a unified base model ${\bm{\theta}}_{0}$. For some modules in the network, differences between the individual models and the base model (i.e., ${\bm{\theta}}_{i}-{\bm{\theta}}_{0}$’s) are maintained. These differences are treated as experts within the framework. Then, a mixture of experts is applied, where a gating network conditioned on user preferences determines the weighting of these experts. It enables the generation of diverse models by applying different weighted combinations of experts to align with specific user preferences.
@@ 5.2 Training Strategy
This section outlines the training approach for the parameters (denoted ${\bm{\phi}}$) of the network structures introduced in Section 5.1. Specifically, ${\bm{\phi}}$ refers to the parameters of the Hypernetwork in Section 5.1.1, parameters of the preference-conditioned network in Section 5.1.2, or parameters of the base models or low-rank models in Section 5.1.3. Current algorithms follow a similar approach: sample user preferences and optimize the structure parameters to generate networks that align with those preferences. The training objective can be written as:
@@
where $({\bm{x}},{\bm{y}})$’s are training data sampled from dataset $\mathcal{D}$, ${\bm{\ell}}(\cdot,\cdot)$ is the multi-objective loss function and $\tilde{g}_{\bm{\alpha}}(\cdot)$ is an MOO algorithm that produces a single solution given preference vector ${\bm{\alpha}}$. In general, most optimization algorithms discussed in Sections 2.2 and 4 can be adopted. Below, we summarize the algorithms adopted by the existing methods.
@@
  • Preference-Aware Weighting Methods: As discussed in Section 4.1, these methods incorporate user preferences into optimization. PHN [52] employs the Exact Pareto Optimal (EPO) solver [37] to find Pareto-optimal solutions aligned with preferences. CPMTL [53] uses a constrained approach inspired by PMTL [36].
  • Hypervolume Maximization: As discussed in Section 4.2, hypervolume maximization can optimize both diversity and convergence of a solution set. PHN-HVI [54] leverages hypervolume maximization to optimize the structural parameters ${\bm{\phi}}$. In each iteration, it samples a set of preference vectors and generates solutions based on these vectors. Subsequently, hypervolume maximization is applied to promote both the diversity and convergence of the solutions.
@@ 5.3 Discussions
Selecting the appropriate structural framework and optimization algorithms is important for effectively learning an infinite Pareto set. Hypernetwork-based approaches offer significant flexibility by generating a diverse set of models tailored to varying user preferences. However, their main drawback is size: the output dimension matches the target network’s parameters, making hypernetworks significantly larger than base models. This limits their applicability for large-scale networks or resource-constrained scenarios.
In contrast, methods based on preference-conditioned networks are highly parameter-efficient. They introduce a minimal number of additional parameters relative to the base model, making them well-suited for large-scale applications where adding extra parameters is costly. However, their limited parameter space may reduce diversity, hindering their ability to approximate complex PS accurately.
Methods based on model combinations serve as a middle ground between hypernetwork and preference-conditioned network methods. By integrating multiple models, they offer a good approximation of the PS while maintaining scalability for large models. Furthermore, some methods allow for parameter adjustments through the selection of appropriate ranks in low-rank approximations.
Regarding training strategies, most algorithms use linear scalarization for its simplicity and strong performance. However, when aligning solutions with user preferences is important, preference-dependent weighting schemes in Section 4.1 are more suitable. Additionally, to achieve HV-optimal solutions, HV-based optimization criteria in Section 4.2 can also be employed, as they better capture the PF by focusing on the dominated objective space volume.
@@ 6 Theories
This section reviews the convergence analysis for gradient-balancing methods in Section 3.2. This is first examined under the deterministic gradient setting (i.e., full-batch gradient). Subsequently, it is analyzed under the stochastic gradient setting. Note that in non-convex MOO, convergence refers to reaching Pareto stationary (Definition 5).
@@ 6.1.1 Deterministic Gradient
With the use of deterministic gradients, it has shown that under mild conditions, MGDA [6] can converge to a Pareto-stationary point at a convergence rate of $\mathcal{O}(K^{-1/2})$ [97], where $K$ is the number of iterations. This rate is comparable to that of single-objective optimization [98]. The fundamental idea is to assess the reduction in each individual objective function. When the current solution is not Pareto-stationary, it can be shown that the common descent direction ${\bm{d}}$ identified by MGDA ensures an improvement of at least $\mathcal{O}(\left\|{\bm{d}}\right\|^{2})$ in each objective, which mirrors the situation in single-objective optimization [97]. Therefore, the method used in single-objective optimization can be applied to prove the convergence. Furthermore, other algorithms such as CAGrad [19], Nash-MTL [27], and PCGrad [33] also offer convergence analyses in similar approaches.
TABLE II: Convergence of existing stochastic gradient MOO algorithms. “LS” and “GS” denote the objectives are $L$-smooth and generalized $L$-smooth, respectively. “BG” and “BF” represent the bounded gradient and bounded function value assumptions, respectively. “Complexity” denotes the sample complexity achieving $\epsilon$-accurate Pareto stationary point. “Bounded CA” denotes the bounded conflict-avoidant distance.
Method Batch Size Assumption Complexity Bounded CA
@@ 6.1.2 Stochastic Gradient
Liu and Vicente [20] are the first to analyze the convergence of gradient-based MOO algorithms with stochastic gradients. They propose a stochastic version of MGDA [6] that computes a common descent direction using stochastic gradients for all objectives and prove convergence to Pareto-optimal solutions under the assumption of convex objective functions. However, due to the inherent bias of common descent direction ${\bm{d}}$ introduced by stochastic gradient estimations, their analysis requires the use of an increasing batch size that grows linearly with the number of iterations to ensure convergence. This requirement can be impractical in real-world applications, as large batch sizes lead to increased computational costs and memory usage.
To overcome this limitation, Zhou et al. [21] propose Correlation-Reduced Stochastic Multi-Objective Gradient Manipulation (CR-MOGM). It addresses the bias in the common descent direction by introducing a smoothing technique on weight coefficient ${\bm{\lambda}}$. In CR-MOGM, ${\bm{\lambda}}^{(k)}$ at iteration $k$ is updated using a moving average:
@@
where $\gamma\in[0,1]$ is a smoothing factor, $\hat{{\bm{\lambda}}}^{(k)}$ is the weight vector obtained by the MOO solver (such as MGDA) at iteration $k$ using the current stochastic gradients, and ${\bm{\lambda}}^{(k-1)}$ is the smoothed weight from the previous iteration. Smoothing reduces the variance in the weight, leading to a more stable and reliable common descent direction. Xu et al. [23] further analyze the convergence by updating the objective weights every $\tau$ iterations instead of at each iteration, and achieve the same convergence rate ${\mathcal{O}}(\epsilon^{-2})$ as CR-MOGM. They also match the convergence rate in stochastic single-objective optimization [99].
@@
$\hat{{\bm{g}}}^{(k+1)}_{i}=\prod_{L_{i}}\left(\hat{{\bm{g}}}^{(k)}_{i}-\gamma\left(\hat{{\bm{g}}}^{(k)}_{i}-{\bm{g}}^{(k)}_{i}\right)\right),$ (42)
where $\prod_{L_{i}}$ is the projection to the set $\{{\bm{g}}\in\mathbb{R}^{d}~{}~{}|~{}~{}\left\|{\bm{g}}\right\|\leq L_{i}\}$, and $L_{i}$ is the Lipschitz constant of $f_{i}({\bm{\theta}})$. However, the analysis requires the number of iterations $K$ to be sufficiently large (in the order of ${\mathcal{O}}(m^{10})$). This high dependency on $m$ makes MoCo less practical for problems involving many objectives. Additionally, the convergence analysis of CR-MOGM and MoCo relies on the assumption that the objective functions have bounded values.
To address these limitations, Chen et al. [24] propose the Multi-objective gradient with double sampling algorithm (MoDo). MoDo mitigates bias in stochastic gradient-based MOO methods without assuming bounded function values. It introduces a double sampling technique to obtain unbiased estimates of the gradient products needed for updating the weight coefficients. Specifically, at each iteration, MoDo updates the weight vector ${\bm{\lambda}}^{(k)}$ using gradients computed on two independent mini-batches:
@@
where $\eta$ is the step size, $\prod_{\Delta_{m-1}}$ is the projection onto the simplex $\Delta_{m-1}$, and ${\bm{G}}^{(k)}({{\bm{z}}^{(k)}_{1}})$, ${\bm{G}}^{(k)}({{\bm{z}}^{(k)}_{2}})$ are stochastic gradients evaluated on two independent mini-batches ${\bm{z}}_{1}^{(k)}$ and ${\bm{z}}_{2}^{(k)}$. By reducing the bias in estimating the common descent direction, MoDo achieves convergence guarantees without the bounded function value assumption. Additionally, it guarantees a bounded conflict-avoidant (CA) distance, which is the distance between the estimated update direction and the CA direction defined in problem (12).
Similarly, Xiao et al. [25] propose the Stochastic Direction-oriented Multi-objective Gradient descent (SDMGrad) algorithm, which introduces a new direction-oriented multi-objective formulation by regularizing the common descent direction within a neighborhood of a target direction (such as the average gradient of all objectives). In SDMGrad, the weight vector ${\bm{\lambda}}^{(k)}$ is updated as:
@@
where $\eta$ is the step size, $\gamma$ is a regularization factor controlling the proximity to the target direction ${\bm{g}}_{0}({\bm{z}}_{2}^{(k)})$. By incorporating this regularization, SDMGrad effectively balances optimization across objectives while guiding the overall descent direction. Zhang et al. [26] further consider the convergence under generalized $L$-smoothness [100] and without bounded gradient assumption. A summary of the convergence results of existing stochastic gradient MOO algorithms is provided in Table II.